A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

13 September 2023 Version 1.05

Title word cross-reference

(p, p, 3) [261]. 0 < p < 1 [207]. 2 [44]. 7 [50]. 8 [50]. ∗ [260]. b [101]. C∗ [1, 79]. C^{∗} [3, 119, 113, 262, 23]. C^{∗} [183]. f [19]. G_{2} [84]. GL_{2} × GU(1) [249]. H [54, 34]. K [70, 215, 171]. L [249, 18, 131]. L^{p} [173]. H^{n} [90]. R^{d} [256]. R^{n} [208]. Z_{p}^{r}(1) [198]. C^{k} [219]. O [28]. ⟨⇕ [233]. ∫↕_{n} [116, 179]. GL_{4} [158]. GL_{n} × GL_{n} \ GL_{2n} [177]. N [231]. n − k [200]. n^{2}α [196]. p [249, 9, 216, 199]. P^{n} [112]. π_{1} [51, 61]. SL(n) [86, 72]. Symp(CP^{2}♯5C̄P^{2}, ω) [259]. × [200].

-adic [249, 216, 199]. -algebras [23, 260, 79, 3, 183, 119, 113, 1]. -categories [54]. -colored [231]. -dimensional [50]. -divergence [19]. -functions [249, 18, 131]. -graphs [171]. -inner [157]. -invariant [34]. -isogenies [248]. -positivity [215]. -regularity [173]. -Semistability [51, 61]. -structures [84]. -symmetry [200]. -symplectic [101]. -theory [70]. -tori [44].

1324- [215].

2143-avoiding [215]. 2D [245].

3-manifolds [139].

4-manifolds [162, 163].

abelian [151, 114]. Absorption [79, 1].

Action [37, 146]. Actions [51, 259, 163, 114]. addable [26]. Addendum [179]. Addition [19]. adic [249, 216, 199]. adjoint [53]. admissible [220, 134]. Albert [106]. Alexander [188]. Algebra [88, 49, 214]. Algebras [52, 62, 78, 63, 8, 42, 221, 59, 106, 263, 172, 3, 23, 258, 128, 183, 236, 135, 119, 113, 117, 218, 1, 260, 105, 262, 79]. Algorithms [214]. Almost [22, 39]. along [186, 155]. Alternating [93]. Analogue [83]. Analytic [75, 257, 241]. anisotropic [193]. Anisotropically [41]. Annelidan [45]. Application [99, 182]. Applications [96, 40, 29, 115]. approach [206, 136]. approximants [159]. Approximation [208, 183]. arbitrary [156]. Arising [8]. arithmetic [194, 216, 158]. Aronszajn [246]. Arthur [229]. Artin [118, 165]. Ascending [61]. Associated [89, 2, 240, 113, 124, 166]. Automata [66]. Automatic [17]. Automaticity [66]. automorphic [147]. Automorphisms [203, 145]. avoiding [215]. Axiomatics [32]. Axioms [122].

B [218]. Back [13, 25, 36, 47, 57, 68, 81, 95, 108, 121, 130, 142, 153, 168, 181, 191, 202, 211, 212, 227, 239, 252, 266]. Bakry [6]. balanced [139]. Banach [128]. Bastiani [219]. Beilinson [92]. Bergman [173]. between [213, 234, 207]. Beyond [43]. Bézout [45]. Bipartite [64, 74]. biquadratic [244]. Blanchfield [188]. blocks [116, 179]. BMO [207]. bodies [157]. Bombieri [40]. Bound [22, 11]. Boundaries [103]. Boundary [11, 255]. Bounded [61]. Boundedness [97]. Bounding [104, 127]. Bounds [240, 214]. braids [123]. Branch [37]. Bratteli [69]. Bridge [26]. Bridge-addable [26]. buildings [203]. Bundles [44, 112].

C [175]. Calabi [21, 89]. calculus [219].

Calderón [55]. Cancellation [31]. Carpenter [60]. case [249, 114, 207]. categories [54, 185]. Categorification [116, 179]. Category [69, 28]. Cauchy [126]. CCR [63]. Cellular [66]. Centre [9]. Centre-by-metabelian [9]. certain [185, 150, 118, 165]. characteristic [250]. characteristics [224]. characterization [264]. Characters [92]. Chen [161]. Cherednik [218]. Chern [204]. Chord [96]. Chow [21]. circle [259]. circles [143, 145]. CJM [12, 13, 24, 25, 35, 36, 46, 47, 56, 57, 67, 68, 80, 81, 94, 95, 107, 108, 120, 121, 129, 130, 141, 142, 152, 153, 167, 168, 180, 181, 190, 191, 201, 202, 209, 210, 211, 212, 226, 227, 238, 239, 251, 252, 265, 266]. class [131]. Classes [26, 189, 244, 197, 187, 208]. Classical [5, 263]. Classification [88, 50]. Classifying [105]. Club [246]. cluster [258]. CM [240, 20]. Co [54]. Co- [54]. cocompact [51]. coefficients [232, 135, 147]. Coercive [72]. Cohomological [220]. cohomologies [112]. Cohomology [54, 154]. Coisotropic [101]. Coleman [104]. colored [231]. colorings [223]. Combinatorial [70]. Combinatorics [189, 205, 62, 229, 218]. Commutator [55]. commutators [138]. Commuting [49]. compactifications [149]. compactness [138]. Comparison [11, 217]. complex [237, 109]. complexes [242]. complexity [148]. compositum [132]. Computing [234]. Concordance [58, 188]. conditions [255, 163]. cones [136, 149]. configurations [193]. conformal [187]. congruences [224]. Conjecture [27, 200, 198, 118, 165, 161]. Connected [82]. connections [185]. connexe [115]. constant [198]. constants [126]. Constellations [74]. construction [115]. continuation [213]. Contravariant [86]. Convergence [26, 146, 136]. Convex [72, 229, 232, 137]. Convolution [104]. cores [113]. correlations [196]. Correspondence [3, 23]. Corrigendum [23, 166]. Cost [16]. Counterpart [15]. Cover [12, 13, 24, 25, 35, 36, 46, 47, 56, 57, 67, 68, 80, 81, 94, 95, 107, 108, 120, 121, 129, 130, 141, 142, 152, 153, 167, 168, 180, 181, 190, 191, 201, 202, 209, 210, 211, 212, 226, 227, 238, 239, 251, 252, 265, 266]. coverings [220]. Covolumes [90]. Coxeter [4, 98]. creep [144]. Crossed [3, 23]. crystal [218]. cube [242]. Curvature [6, 11]. Curves [20, 194, 240, 248]. Cuspidal [88]. Cyclic [110, 132]. Cyclicity [194]. Cynk [21].

data [115]. defined [126]. Definitions [96]. Deformation [186]. Degenerating [187]. degree [132]. Degrees [74]. denominators [205]. densely [126]. Densities [150, 197]. Density [18, 193]. Depth [61]. Derivatives [160]. Diagrams [69, 98, 189, 205, 162]. dichotomy [126]. difference [134]. Differential [97, 219, 134, 214]. Dimension [113]. Dimensional [90, 172, 198, 89, 76, 50]. dimensions [205]. Diophantine [261]. Dirac [186]. Direct [27]. directional [176]. Dirichlet [131]. Discrete [3, 23]. discriminants [161]. Discs [75]. distribution [240]. Distributive [45]. divergence [19]. Dividing [20]. Divisibility [151]. Division [59]. Domains [82, 41, 173, 156, 247]. Données [115]. double [220, 125]. Doubles [58]. dreams [195]. drift [124]. Drinfeld [241]. duality [235]. d’un [115]. d’une [115]. Dunkl [184]. dyadic [256]. dynamical [114].

Effective [241]. Eigencurve [5]. Eigenvalue [41, 111]. eigenvalues [160]. Eisenstein [243, 8]. Elements [92]. elliptic [194, 240, 248]. Émery [6]. Endoscopic [115]. endoscopiques [115]. endpoint [124]. energy [225]. Enriques [2, 166]. Entropy [91]. epsilon [198]. equation [245, 169]. equations [134, 261]. Equisingular [89]. equivalence [189, 171]. Equivariant [73, 154]. Erdos [71]. Erratum [79, 165].

Estimates [102, 124]. étale [105, 135]. Euclidean [124]. Euler [245, 224]. Evolution [85]. Exact [230]. exceptional [203]. Existence [245, 85]. exotic [143]. Expanding [41]. exponent [223]. exponential [126]. Extension [100, 219]. Extensions [78, 61, 244]. Extremal [137].

Factorization [109]. factorizations [222]. Families [104, 174]. Family [89, 144]. Fiber [10]. Fibrations [237]. fibred [125]. Fields [110, 102, 151, 132, 261, 248]. Finite [16, 49, 33, 102, 172, 163]. finite-dimensional [172]. Finiteness [192, 128]. Finsler [144]. First [27, 111]. Flag [4]. Flat [41]. Flow [37]. flows [204, 155]. Fock [257]. forcing [122]. Form [110]. Forms [49, 162]. formula [229, 126, 140, 169]. Formulas [74]. Fourier [147, 39]. fractional [111, 196]. Free [114, 103, 59, 91, 33, 71]. French [115]. Frobenius [231, 240]. Front [12, 24, 35, 46, 56, 67, 80, 94, 107, 120, 129, 141, 152, 167, 180, 190, 201, 209, 210, 226, 238, 251, 265]. Function [53]. Functions [6, 72, 228, 249, 257, 18, 131]. Functoriality [30]. Fundamental [158, 259]. Further [157].

Galois [23, 3, 244]. Gauss [83]. GCR [63]. Gelfand [263, 260]. general [256]. Generalised [17]. Generalization [29, 242, 199]. Generalizations [96, 85]. Generalized [92, 48, 231, 237]. Generated [97]. generator [262]. generators [105]. Gentle [62, 78]. Genus [64]. geometric [137]. Geometry [99, 101, 11, 189, 205, 237, 149]. Gleason [75]. global [249]. Graev [263]. Graph [64, 197]. Graphs [6, 146, 171]. Grassmann [205, 254]. Grassmannian [258, 149]. Grassmannians [254]. Green [53]. grids [256]. Group [51, 84, 59, 259, 206, 163, 114, 115, 199]. group-theoretic [199]. group-theoretical [206]. groupe [115]. groupoid [135]. Groups [4, 3, 104, 40, 90, 9, 263, 223, 147, 113, 127, 109, 222, 216, 224, 192, 118, 165, 23]. Growth [9].

Half [15]. Half-integral [15]. Hamilton [169]. Hamiltonian [186]. Hardy [207]. Harmonic [82, 234]. Heat [97]. Hecke [263, 18]. Heisenberg [44]. Hessian [2, 166]. Higher [40, 90, 55, 198]. HNN [61]. Hodge [89]. Hölder [208]. Holomorphic [37, 164]. Homology [9, 170, 123]. Homotopy [188]. Hopf [172]. Hulek [21]. hypergraphs [137]. Hyperkähler [99].

Ideal [75, 90]. identities [231]. II [205, 59, 91]. immanants [215]. Inclusions [79, 1]. incompressible [245]. indefinite [111]. independent [134]. Index [96, 164]. Inequalities [77, 157]. inequality [176]. Infinite [76, 123]. Infinite-dimensional [76]. Infinity [51, 61]. injective [235]. inner [157]. Inputs [34]. integrable [185]. integral [15]. integrals [177]. Interpolation [207, 249]. intersection [162]. intertwining [216]. Intrinsic [7]. Invariant [72, 66, 37, 140, 34]. Invariants [223, 58, 20, 220, 127, 139]. inverse [135]. Invertibility [221]. Involution [195, 236]. Involutions [76]. Irredundant [42]. isogenies [248]. isomorphism [118, 165]. isomorphisms [87]. Isospectrality [14]. Issue [12, 13, 24, 25, 35, 36, 46, 47, 56, 57, 67, 68, 80, 81, 94, 95, 107, 108, 120, 121, 129, 130, 141, 142, 152, 153, 167, 168, 180, 181, 190, 191, 201, 202, 209, 210, 211, 212, 226, 227, 238, 239, 251, 252, 265, 266]. Iwahori [52]. Iwasawa [127, 29].

Jacobi [169]. Jacobians [253]. Jiang [79, 1]. Jordan [8].

Kac [117]. Kadison [60]. Kähler [185]. Kazhdan [215]. kernels [184]. Khovanov [170, 123]. knotted [143]. Kolitsch [231]. Köthe [98]. Kuperberg [139].

laced [147]. Landsberg [200]. Langlands [115, 30]. Laplacian [124]. Laplacians [184, 53]. Large [42, 131, 84]. Lattice [88, 41]. Lattices [65]. Laugwitz [200]. Lemma [158]. length [192]. Lens [14]. Level [18]. Liberation [91]. Lidskii [160]. Lie [88, 28]. line [264]. Linear [66, 134, 158, 174, 192]. Linearity [87]. linearly [134]. Links [58]. Liouville [255]. Lipschitz [33]. Lipschitz-free [33]. LMO [140]. Local [26, 249, 155, 199]. local-global [249]. locally [178]. Loewner [247]. log [232]. log-convex [232]. loop [189, 205]. Loops [259]. Low [18]. Low-lying [18]. Lower [22, 11, 112]. Lucas [83]. Lusztig [215]. lying [18]. LYZ [86].

Mallows [133]. Manifolds [11, 185, 213, 162, 186, 234, 163, 139]. many [145]. Maps [74, 213, 182, 234, 113, 247]. Maranda [235]. Marginals [16]. Martingale [77, 207]. Masures [32]. matchings [133]. Matrices [86, 172]. Matrix [91, 113]. matroids [189]. matter [12, 13, 24, 25, 35, 36, 46, 47, 56, 57, 67, 68, 80, 81, 94, 95, 107, 108, 120, 121, 129, 130, 141, 142, 152, 153, 167, 168, 180, 181, 190, 191, 201, 202, 209, 210, 211, 212, 226, 227, 238, 239, 251, 252, 265, 266]. Maximal [75, 77, 55]. maximum [155]. McKay [172]. Measurable [60]. Measure [82]. Measures [19, 66, 39, 264]. median [242]. Menchov [85]. metabelian [223, 9]. Metric [33]. metrics [144, 243]. Metrized [44]. minimal [147]. minimization [111]. Minkowski [200]. Minus [15]. mirror [125]. Mixed [4, 236, 89]. Möbius [145]. Models [30]. Modular [90, 116, 179]. module [244]. Modules [54, 88, 73, 44, 10, 235, 241, 230]. modulo [194]. moment [264]. moments [253]. monodromy [199]. Moody [117]. Morita [171, 31]. Morse [204]. Moser [71]. Motives [73]. Moves [171]. multicolored [123]. multiplication [228]. multiplicative [225]. Multiplicity [250]. multipliers [178]. multisigned [224].

Naimark [260]. Naturally [50]. Nearly [84]. necklaces [205, 254]. Neumann [49, 43]. Nevanlinna [221]. Newforms [15]. Newton [149]. next [147]. next-to-minimal [147]. no [75]. Non [37, 51, 240, 126]. non-CM [240]. Non-cocompact [51]. non-densely [126]. Non-invariant [37]. nonautonomous [126]. Noncommutative [77, 206, 207]. noncompact [186]. nonnegative [254]. Nontame [204]. normal [148, 114]. norms [200]. Number [110, 151, 233, 134, 261, 150]. numbers [148].

odd [204, 132]. Okounkov [149]. One [18, 89, 250]. One-dimensional [89]. One-Level [18]. Operator [42, 55, 230, 247]. Operators [97, 49, 85, 146, 255, 186, 124, 216]. opposition [203]. Optimal [159]. Optimisation [41]. Optimization [19]. Orbifold [14]. orbifolds [154]. Orbital [91, 177]. orbits [186]. Order [38, 55, 87]. Order-isomorphisms [87]. Ordered [69, 137, 260]. Orlicz [19]. orthogonal [132, 159]. Oscillating [178].

p [157, 248]. pairings [188]. pairs [228]. Parallel [84, 157]. parametric [174]. Parisi [169]. partitions [231]. Parts [75, 196]. paths [156]. PDE [217]. peeling [182]. percolation [182]. Periodic [39]. Periplectic [28]. permutations [133]. perturbation [255]. Perverse [4]. Peter [52]. Picard [20]. pipe [195]. planar [182]. Plane [37, 156]. planes [125, 149]. point [193]. Points [5, 41]. polyharmonic [213]. Polyhedra [90]. Polyhedral [53]. Polynomiality [222]. Polynomials [17, 70, 34, 232, 206, 188, 159]. Polytopes [86, 93, 22, 189, 229, 136]. Posets [38]. positive [250]. positivity [215]. Potential [184]. potentials [258]. power [223, 244]. Prescribed [74]. preserving [171]. Prikry [122]. Prime [34, 223, 132, 150]. Primes [20, 194]. Primitive [28]. principles [217, 155]. Pro [9]. Pro- [9]. probability [206]. Problem [19, 31, 71, 111, 174, 98]. problems [257, 144, 137, 217, 109, 126]. Process [91]. Products [27, 3, 23, 76, 65, 240, 128]. progressions [194]. projection [173]. Projections [65]. projective [114]. Proof [200, 158]. properties [213, 157]. Property [82, 100, 246, 216]. pure [235]. pure-injective [235].

Quadratic [18, 248]. Quadrics [7]. quantization [186]. Quantum [44, 175]. Quartic [18, 2, 166]. quasi [242]. quasi-median [242]. questions [174]. quiver [254]. Quivers [258]. Quotient [221, 173].

races [150]. Rademacher [85]. radial [184]. Radius [102]. Ramification [5]. Random [64, 257, 136]. Rank [163, 40, 114, 262]. Rankin [104]. Rankin-Selberg [104]. Rapoport [48]. rational [218]. real [264]. recurrence [174]. Reduced [3, 23, 128]. réductif [115]. Reductive [50, 228, 115]. Reflection [103, 246, 109, 222]. regions [170]. regularity [173]. related [106, 115, 174]. Relation [91]. relations [119]. Relative [73]. representation [263]. Representations [30, 147, 113, 116, 179, 218]. represented [259]. Repulsive [16]. Restriction [102]. Resultants [161]. results [148]. revisit [138]. revisited [242]. ribbon [188]. Ricci [155, 11]. Riemannian [213, 234]. rigid [241]. Ring [128, 21]. Ring-theoretic [128]. Rings [45, 10, 59, 106, 236, 135]. Robin [43, 217]. root [220]. roots [232]. rotation [119]. Rozansky [123].

same [132]. SCFTs [164]. Schmidt [34]. Schreieder [21]. Schrödinger [97, 85]. Sector [83]. Seeley [219]. Seeley-extension-theorem [219]. Selberg [104]. Selector [60]. Self [53, 113]. Self-adjoint [53]. self-similar [113]. Selmer [104, 127, 224]. semigroup [135]. Semigroups [97]. Semiregular [93]. Semistability [51, 61]. semitoric [237]. Sequences [17, 187, 174]. Series [8]. Set [71]. Sets [42, 100]. several [159]. sextic [125]. Sharp [124]. sheaf [135]. Sheaves [4]. sheets [245]. Shortest [156]. Sigma [122]. signature [261]. similar [113]. Simple [88, 182, 3, 23, 74, 65, 183]. Simplicial [22]. Simply [82, 147]. simply-laced [147]. Simultaneous [49]. Singular [37, 255]. site [182]. skew [135]. Skolem [174]. Slice [58]. Slice-torus [58]. Slices [99]. Smale [204]. small [164]. smooth [143]. soluble [192]. solutions [134]. Some [148, 155, 48, 124, 174, 192]. Soulé [92]. Space [15, 76, 124, 169]. Spaces [143, 14, 33, 75, 48, 50, 257, 178, 207, 105]. sparse [197]. Special [242, 10, 246]. spectral [264]. Spectrum [28, 187]. Sphere [102]. Spherical [99, 228, 203]. splicing [140]. square [244]. Stability [26, 119]. stable [133]. stacks [220]. stairs [223]. stationary [246, 245]. Steinberg [63]. Steiner [111]. Steklov [43, 187]. Strongly [39]. structure [244, 117, 118, 165]. Structures [89, 143, 230, 84]. Sturm [255]. subgroups [151]. subhomogeneous [262]. Submanifolds [101]. submaximal [114]. subsets [225]. Sum [71]. Sum-free [71]. Super [72]. Super-Coercive [72]. Superalgebra [28]. Surface [2, 166].

Surfaces [53, 103, 151, 145]. sutured [139]. Swan [29]. symmetric [178]. Symmetry [84, 111, 200]. symplectic [101, 160]. Szego [43].

Tannakian [185]. terms [263]. ternary [261]. their [65, 224]. Theorem [40, 29, 235, 219, 160, 247, 199, 60, 85, 83]. Theorems [22, 193, 250, 260]. theoretic [128, 199]. theoretical [206]. Theory [29, 148, 255, 204, 70]. thin [170]. three [150]. three-way [150]. Threefolds [125, 89]. Threshold [221]. Tight [26]. Topology [75]. Tori [41, 44]. toric [229]. Torsion [170, 151]. torsional [144]. Torsions [162]. torsors [106]. torus [58, 154, 186]. Totally [254]. Trace [110, 229]. traces [240]. Tracial [183]. Transformability [39]. Transforms [97, 77]. Tree [197, 246]. Trees [65]. Triangular [49]. triple [196]. trisection [162]. trivializations [241]. Tropical [149, 253]. twisted [188]. twists [198]. Two [257, 264]. Two-moment [264]. Type [88, 228, 160, 218, 203, 175].

ultrafilter [233]. Unicorn [200]. Uniform [90]. Unique [213]. Unital [79, 1]. Unitary [218]. Universal [100, 93]. unramified [198]. Upper [49, 22, 214].

Valuation [59]. Valuations [86, 72]. values [131]. vanishing [112]. variables [159]. Variation [126, 89]. Varieties [4, 21, 229, 114, 225]. Vector [44, 76]. via [189, 255, 249, 43, 116, 179]. Vinogradov [40]. volume [12, 13, 24, 25, 35, 36, 46, 47, 56, 57, 67, 68, 80, 81, 94, 95, 107, 108, 120, 121, 129, 130, 141, 142, 152, 153, 167, 168, 180, 181, 190, 191, 201, 202, 209, 210, 211, 212, 226, 227, 238, 239, 251, 252, 265, 266]. vortex [245].

Wasserstein [169]. Wave [85]. Way [43, 150]. Weak [38, 136]. webs [175]. Weight [15, 111]. Weighted [11]. Weil [204]. Weyl [52, 38]. which [259]. Whitehead [58]. Whitney [176]. Wick [206]. Willmore [103]. Wilson [189, 205]. Witt [88, 236].

Yau [21, 89]. yield [214]. Yui [161].

Zassenhaus [27]. Zero [102]. Zeroes [34]. Zeros [18, 131]. Zink [48]. Zygmund [208].

References

Osaka:2018:JAI

[1] Hiroyuki Osaka and Tamotsu Teruya. The Jiang–Su absorption for inclusions of unital C∗-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(2):400–??, April 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [79].

Shimada:2019:ESA

[2] Ichiro Shimada. On an Enriques surface associated with a quartic Hessian surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(1):213–246, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-an-enriques-surface-
associated-with-a-quartic-hessian-
surface/7F4ED300013922C9D14F240FEA1B5DC4.

See corrigendum [166].

Cameron:2019:GCR

[3] Jan Cameron and Roger R. Smith. A Galois correspondence for reduced crossed products of simple C^{∗}-algebras by discrete groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(5):1103– 1125, October 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/galois-correspondence-for-
reduced-crossed-products-of-simple-
textcast-algebras-by-discrete-groups/
EE5EC67CFB2D19038582AB6903E93502.

See corrigendum [23].

Achar:2020:MPS

[4] Pramod N. Achar, Simon Riche, and Cristian Vay. Mixed perverse sheaves on flag varieties for Coxeter groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(1):1–55, February 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/mixed-perverse-sheaves-
on-flag-varieties-for-coxeter-groups/
84B6A009DDB994CB4CC2B8E326EBC36B.

Betina:2020:REC

[5] Adel Betina. Ramification of the eigencurve at classical RM points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (1):57–88, February 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/ramification-of-the-eigencurve-
at-classical-rm-points/66D74ADE59625A33A7D78151A63D8646.

Cushing:2020:BEC

[6] David Cushing, Shiping Liu, and Norbert Peyerimhoff. Bakry–Émery curvature functions on graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (1):89–143, February 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/bakryemery-curvature-functions-
on-graphs/2CE894DF2E796AABD5AA0920C415988E.

Fahrner:2020:IQ

[7] Anne Fahrner and Jürgen Hausen. On intrinsic quadrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(1):145– 181, February 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-intrinsic-quadrics/43A914E95CADE648B1628EAF7824992A.

Hanzer:2020:ESA

[8] Marcela Hanzer and Gordan Savin. Eisenstein series arising from Jordan algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(1):183– 201, February 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/eisenstein-series-arising-
from-jordan-algebras/AE032F11A3717F7175A486BBE537549F.

Kochloukova:2020:GHC

[9] Dessislava H. Kochloukova and Aline G. S. Pinto. Growth of homology of centre-by-metabelian pro-p groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (1):203–224, February 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/growth-of-homology-of-centrebymetabelian-
pro-p-groups/7DD917C46AF3557FDCAD118776C42848.

Miranda-Neto:2020:SFR

[10] Cleto B. Miranda-Neto. On special fiber rings of modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (1):225–242, February 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-special-fiber-rings-
of-modules/21F51D45EE470B83EA8012E24D5BA8E3.

Sakurai:2020:CGM

[11] Yohei Sakurai. Comparison geometry of manifolds with boundary under a lower weighted Ricci curvature bound. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (1):243–280, February 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/comparison-geometry-of-
manifolds-with-boundary-under-a-
lower-weighted-ricci-curvature-
bound/0B17AB82188C0EED67D91FD3786DE316.

Anonymous:2020:CVIa

[12] Anonymous. CJM volume 72 issue 1 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (1):f1–f2, February 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-72-issue-1-cover-
and-front-matter/2CD03D94E56F18F8D0EE72EAD5F8CBAE.

Anonymous:2020:CVIb

[13] Anonymous. CJM volume 72 issue 1 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (1):b1–b2, February 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-72-issue-1-cover-
and-back-matter/6E23519B4B168B550809C74B0E38E15F.

Bari:2020:IOL

[14] Naveed S. Bari and Eugenie Hunsicker. Isospectrality for orbifold lens spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):281–325, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/isospectrality-for-orbifold-
lens-spaces/20B041963BE8C514B65E642E76F923C0.

Baruch:2020:NHI

[15] Ehud Moshe Baruch and Soma Purkait. Newforms of half-integral weight: The minus space counterpart. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (2):326–372, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/newforms-of-halfintegral-
                                        weight-the-minus-space-counterpart/
                                        41E25686F086390C7CE8D31099DF4D0A.

Bindini:2020:MFR

[16] Ugo Bindini. Marginals with finite repulsive cost. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):373–391, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/marginals-
                                        with-finite-repulsive-cost/6D5C58646A2F5B90D7192C4F09C71676.

Byszewski:2020:ASG

[17] Jakub Byszewski and Jakub Konieczny. Automatic sequences and generalised polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):392– 426, April 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/automatic-sequences-and-
                                        generalised-polynomials/89086F399E83E464032E523A6DD2BEAB.

Gao:2020:OLD

[18] Peng Gao and Liangyi Zhao. Onelevel density of low-lying zeros of quadratic and quartic Hecke Lfunctions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):427–454, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/onelevel-
density-of-lowlying-zeros-of-quadratic-
and-quartic-hecke-l-functions/AA8B6192E14B218D85780226477A0D3F.

Hou:2020:OAM

[19] Shaoxiong Hou and Deping Ye. Orlicz addition for measures and an optimization problem for the f-divergence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):455–479, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/orlicz-addition-for-measures-
and-an-optimization-problem-for-
the-f-divergence/B0A592E0D16EDB30F4F5AF83D2153DE8.

Kilicer:2020:PDI

[20] Pinar Kiliçer, Elisa Lorenzo García, and Marco Streng. Primes dividing invariants of CM Picard curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):480–504, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/primes-dividing-invariants-
of-cm-picard-curves/ECFF74081A06CABFD749070B0A916FCA.

Laterveer:2020:CRC

[21] Robert Laterveer and Charles Vial. On the Chow ring of Cynk–Hulek Calabi– Yau varieties and Schreieder varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):505–536, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-chow-ring-of-cynkhulek-
calabiyau-varieties-and-schreieder-
varieties/5BA7C82B78702B4DEB3E73610A16A7E2.

Nevo:2020:ASP

[22] Eran Nevo, Guillermo Pineda-Villavicencio, Julien Ugon, and David Yost. Almost simplicial polytopes: The lower and upper bound theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (2):537–556, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/almost-simplicial-polytopes-
the-lower-and-upper-bound-theorems/
3605653093E68DA186538F66BFB1C6CC.

Cameron:2020:CGC

[23] Jan Cameron and Roger R. Smith. Corrigendum to: A Galois Correspondence for Reduced Crossed Products of Simple C^{∗}-algebras by Discrete Groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):557–562, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/corrigendum-to-a-galois-
correspondence-for-reduced-crossed-
products-of-simple-textcast-algebras-
by-discrete-groups/C9B9CAAF1F3BE9677AF9595EE3DD5CC7.

See [3].

Anonymous:2020:CVIc

[24] Anonymous. CJM volume 72 issue 2 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (2):f1–f2, April 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-72-issue-2-cover-
and-front-matter/098BA0C9924B5CF5A79990447A0F22C9.

Anonymous:2020:CVId

[25] Anonymous. CJM volume 72 issue 2 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (2):b1–b2, April 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-72-issue-2-cover-
and-back-matter/B2965DF5BBF2A404AA05DC7CAA037CCC.

Chapuy:2020:LCS

[26] G. Chapuy and G. Perarnau. Local convergence and stability of tight bridge-addable classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3): 563–601, June 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/local-convergence-and-stability-
of-tight-bridgeaddable-classes/
009925B8867FB470285769BBB375C8E7.

Bachle:2020:FZC

[27] Andreas Bächle, Wolfgang Kimmerle, and Mariano Serrano. On the first Zassenhaus conjecture and direct products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3):602–624, June 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-the-first-zassenhaus-
                                        conjecture-and-direct-products/
                                        C8D89692709B94CF27DF46E51592D6FD.

Chen:2020:PSC

[28] Chih-Whi Chen and Kevin Coulembier. The primitive spectrum and category O for the periplectic Lie superalgebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3):625–655, June 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/primitive-
                                        spectrum-and-category-mathcalo-
                                        for-the-periplectic-lie-superalgebra/
                                        E43BE718ED452131331C4798929A5416.

Nickel:2020:GTS

[29] Andreas Nickel. A generalization of a theorem of Swan with applications to Iwasawa theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3): 656–675, June 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/generalization-of-a-theorem-
                                        of-swan-with-applications-to-iwasawa-
                                        theory/500808F6E13A6AFDA6A3674728E901E9.

Mitra:2020:MRL

[30] Arnab Mitra and Eitan Sayag. Models of representations and Langlands functoriality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3):676–707, June 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/models-of-
representations-and-langlands-functoriality/
CFF8FB070D2A967033921AEEEEA89410.

Lu:2020:MCP

[31] D.-M. Lu, Q.-S. Wu, and J. J. Zhang. A Morita cancellation problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3): 708–731, June 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/morita-cancellation-problem/
1474505A3639AC3C19372F040D0C0EC7.

Hebert:2020:NAM

[32] Auguste Hébert. A new axiomatics for masures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3):732– 773, June 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/new-axiomatics-for-masures/
F2FB2767CE7D811B228BA9F4B5D30C2F.

Dilworth:2020:LFS

[33] Stephen J. Dilworth, Denka Kutzarova, and Mikhail I. Ostrovskii. Lipschitzfree spaces on finite metric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3):774–804, June 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/lipschitzfree-spaces-on-
                                        finite-metric-spaces/CEA5CE450016D0FCD99EA5B64F333FCA.

Xiao:2020:ZPP

[34] Stanley Yao Xiao and Shuntaro Yamagishi. Zeroes of polynomials with prime inputs and Schmidt’s h-invariant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(3):805–833, June 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/zeroes-of-
                                        polynomials-with-prime-inputs-and-
                                        schmidts-h-invariant/1EF42E517E5343654EC3346F606FE37A.

Anonymous:2020:CVIe

[35] Anonymous. CJM volume 72 issue 3 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (3):f1–f2, June 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-72-issue-3-cover-
                                        and-front-matter/F8B845F983CA0A074C9D8EA4CEF485F9.

Anonymous:2020:CVIf

[36] Anonymous. CJM volume 72 issue 3 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (3):b1–b2, June 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/

canadian-journal-of-mathematics/
article/cjm-volume-72-issue-3-cover-
and-back-matter/C6EFF565A03C481A85F379930C36B253.

Ayuso:2020:APS

[37] P. Fortuny Ayuso and J. Ribón. The action of a plane singular holomorphic flow on a non-invariant branch. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (4):835–866, August 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/action-of-a-plane-singular-
holomorphic-flow-on-a-noninvariant-
branch/12C90BC3B1193929B381DE9059CBE63B.

Gay:2020:WOW

[38] Joël Gay and Vincent Pilaud. The weak order on Weyl posets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (4):867–899, August 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/weak-order-on-weyl-posets/
33AFC4F4F1DE4A94BA13974C71B74A6F.

Strungaru:2020:FTS

[39] Nicolae Strungaru. On the Fourier transformability of strongly almost periodic measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(4):900– 927, August 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-fourier-transformability-

                                        of-strongly-almost-periodic-measures/
                                        2CE9E7A78DAE229B815EB75980416D74.

Jiang:2020:BVT

[40] Yujiao Jiang and Guangshi Lü. The Bombieri–Vinogradov theorem on higher rank groups and its applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (4):928–966, August 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/bombierivinogradov-theorem-
                                        on-higher-rank-groups-and-its-applications/
                                        9AC8E4C30CAC0BDBBE3C66DFF9805F21.

Lagace:2020:EOF

[41] Jean Lagacé. Eigenvalue optimisation on flat tori and lattice points in anisotropically expanding domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (4):967–987, August 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/eigenvalue-optimisation-
                                        on-flat-tori-and-lattice-points-
                                        in-anisotropically-expanding-domains/
                                        351BC9BC029E0FE2B10FFF081C20A50D.

Hida:2020:LIS

[42] Clayton Suguio Hida and Piotr Koszmider. Large irredundant sets in operator algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(4):988–1023, August 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/

article/large-irredundant-sets-
in-operator-algebras/9643E62933C58BA1633BC9F0DDC87ACA.

Freitas:2020:SNB

[43] Pedro Freitas and Richard S. Laugesen. From Steklov to Neumann and beyond, via Robin: The Szegő way. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (4):1024–1043, August 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/from-steklov-to-neumann-
and-beyond-via-robin-the-szego-
way/8FC567395CD94ED07344A5461D91F0DE.

Latremoliere:2020:HMQ

[44] Frédéric Latrémolière. Heisenberg modules over quantum 2-tori are metrized quantum vector bundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(4):1044– 1081, August 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/heisenberg-modules-over-
quantum-2tori-are-metrized-quantum-
vector-bundles/56CF993DD46255437800545189E272CE.

Marks:2020:ADB

[45] Greg Marks and Ryszard Mazurek. On annelidan, distributive, and Bézout rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(4):1082–1110, August 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/

                                        article/on-annelidan-distributive-
                                        and-bezout-rings/71114E3BA94ACC3511CD5099BF86F6E1.

Anonymous:2020:CVIg

[46] Anonymous. CJM volume 72 issue 4 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(4): f1–f2, August 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-72-issue-4-cover-
                                        and-front-matter/94C015D92A7065A35FC697A3E9639E5B.

Anonymous:2020:CVIh

[47] Anonymous. CJM volume 72 issue 4 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (4):b1–b2, August 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-72-issue-4-cover-
                                        and-back-matter/DE18AF45142CB3F245C5480A52DF0099.

Shen:2020:SGR

[48] Xu Shen. On some generalized Rapoport–Zink spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(5):1111– 1187, October 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-some-generalized-rapoportzink-
                                        spaces/B342C5CDB3CB6098F549CA0DA6425F5B.

Charlesworth:2020:SUT

[49] Ian Charlesworth, Ken Dykema, Fedor Sukochev, and Dmitriy Zanin. Simultaneous upper triangular forms for commuting operators in a finite von Neumann algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(5):1188– 1245, October 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/simultaneous-upper-triangular-
forms-for-commuting-operators-in-
a-finite-von-neumann-algebra/483FA9D629E327EC7F419A3F58493F49.

Storm:2020:CDN

[50] Reinier Storm. The classification of 7- and 8-dimensional naturally reductive spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(5):1246– 1274, October 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/classification-of-7-and-
8dimensional-naturally-reductive-
spaces/7264F2621A479860162A3CE3C882B597.

Geoghegan:2020:NCG

[51] Ross Geoghegan, Craig Guilbault, and Michael Mihalik. Non-cocompact group actions and π_{1}-semistability at infinity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (5):1275–1303, October 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/

                                        article/noncocompact-group-actions-
                                        and-unicodestixx1d70b1-semistability-
                                        at-infinity/094659F13DDAD6776F2F19DB2E1FFC08.

Barbasch:2020:PWI

[52] Dan Barbasch and Allen Moy. Peter– Weyl Iwahori algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(5):1304– 1323, October 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/peterweyl-iwahori-algebras/
                                        9721861E5ECB19EAEED0D16905547645.

Kokotov:2020:GFS

[53] Alexey Kokotov and Kelvin Lagota. Green function and self-adjoint Laplacians on polyhedral surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(5): 1324–1351, October 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/green-function-and-selfadjoint-
                                        laplacians-on-polyhedral-surfaces/
                                        84EBE2182B3F52C0A1A1E948E9B12455.

Balodi:2020:CMC

[54] Mamta Balodi, Abhishek Banerjee, and Samarpita Ray. Cohomology of modules over H-categories and co-H-categories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (5):1352–1385, October 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cohomology-of-modules-over-

h-categories-and-co-h-categories/
48E7791BA7D73F273673D5010DC6534B.

Lai:2020:MOH

[55] Xudong Lai. Maximal operator for the higher order Calderón commutator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (5):1386–1422, October 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/maximal-operator-for-the-
higher-order-calderon-commutator/
A1DEC5E45EE50B7D69DD31EF09293376.

Anonymous:2020:CVIi

[56] Anonymous. CJM volume 72 issue 5 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(5): f1–f2, October 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-72-issue-5-cover-
and-front-matter/462D9AE0338D0F02CB0A0CB96D34DFB5.

Anonymous:2020:CVIj

[57] Anonymous. CJM volume 72 issue 5 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (5):b1–b2, October 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-72-issue-5-cover-
and-back-matter/7A41CAB71B5FF22092677243445EADAA.

Cavallo:2020:STC

[58] Alberto Cavallo and Carlo Collari. Slice-torus concordance invariants and Whitehead doubles of links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(6): 1423–1462, December 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/slicetorus-concordance-
invariants-and-whitehead-doubles-
of-links/12FFA1D91D5ED09328D78E9DFF00E381.

Sanchez:2020:FGA

[59] Javier Sánchez. Free group algebras in division rings with valuation II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(6): 1463–1504, December 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/free-group-algebras-in-
division-rings-with-valuation-ii/
267F1A799E2062452AE441DD0F3B06D6.

Bownik:2020:MSK

[60] Marcin Bownik and Marcin Szyszkowski. A measurable selector in Kadison’s Carpenter’s Theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(6):1505– 1528, December 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/measurable-selector-in-
kadisons-carpenters-theorem/26A80B51EB91A0D01053252AA0F91898.

Mihalik:2020:BDA

[61] Michael L. Mihalik. Bounded depth ascending HNN extensions and π_{1}- semistability at infinity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(6):1529– 1550, December 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/bounded-depth-ascending-
hnn-extensions-and-unicodestixx1d70b1-
semistability-at-infinity/914D7612F9B6B965F813FA1D79E2E6BC.

Brustle:2020:CGA

[62] Thomas Brüstle, Guillaume Douville, Kaveh Mousavand, Hugh Thomas, and Emine Yildirim. On the combinatorics of gentle algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(6):1551– 1580, December 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-combinatorics-of-
gentle-algebras/DAC0E0FF011FBAC0FD9E45F473AE2E99.

Clark:2020:GCS

[63] Lisa O. Clark, Benjamin Steinberg, and Daniel W. van Wyk. GCR and CCR Steinberg algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(6): 1581–1606, December 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/gcr-and-ccr-steinberg-algebras/
C5A1246A420BAEDA0CCB716ECC4B560C.

Jing:2020:GRB

[64] Yifan Jing and Bojan Mohar. The genus of a random bipartite graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(6): 1607–1623, December 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/genus-of-a-random-bipartite-
                                        graph/AB463CE386957B88BE99FD1C9D2C42D5.

Radu:2020:NSL

[65] Nicolas Radu. New simple lattices in products of trees and their projections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (6):1624–1690, December 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/new-simple-lattices-in-
                                        products-of-trees-and-their-projections/
                                        999B9869D38B207D69714A9B3A17992D.

Rowland:2020:AIM

[66] Eric Rowland and Reem Yassawi. Automaticity and invariant measures of linear cellular automata. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(6):1691– 1726, December 2020. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/automaticity-and-invariant-
                                        measures-of-linear-cellular-automata/
                                        4756BA06195903E4075BFA34195A406E.

Anonymous:2020:CVIk

[67] Anonymous. CJM volume 72 issue 6 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (6):f1–f2, December 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-72-issue-6-cover-
and-front-matter/54F6117F5E3C94A313A598FAC59F5AFF.

Anonymous:2020:CVIl

[68] Anonymous. CJM volume 72 issue 6 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72 (6):b1–b2, December 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-72-issue-6-cover-
and-back-matter/5126CF8B76BCDF80C65929B897E878A3.

Amini:2021:COB

[69] Massoud Amini, George A. Elliott, and Nasser Golestani. The category of ordered Bratteli diagrams. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (1):1–28, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/category-of-ordered-bratteli-
diagrams/7120B1E22089D9012AF4EC73197A302B.

Monical:2021:PCT

[70] Cara Monical, Oliver Pechenik, and Dominic Searles. Polynomials from combinatorial K-theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(1):29– 62, February 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/polynomials-from-combinatorial-
ktheory/8B6B15CE5E346E81D78EFAAB7CB86380.

Sanders:2021:EMS

[71] Tom Sanders. The Erdős–Moser sum-free set problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(1):63– 107, February 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/erdosmoser-sumfree-set-
problem/6F2D1E5B15357F3FBCFCFDF6E0440A49.

Mussnig:2021:IVS

[72] Fabian Mussnig. SL(n) invariant valuations on super-coercive convex functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(1):108–130, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/textsln-invariant-valuations-
on-supercoercive-convex-functions/
1AAF853AC882CF30E1763EA3134117A3.

Calmes:2021:REM

[73] Baptiste Calmès, Alexander Neshitov, and Kirill Zainoulline. Relative equivariant motives and modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (1):131–159, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/relative-equivariant-motives-
and-modules/7DA22758E93DA78D9559928BEA61B8B6.

Louf:2021:SFC

[74] Baptiste Louf. Simple formulas for constellations and bipartite maps with prescribed degrees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(1):160– 176, February 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/simple-formulas-for-constellations-
and-bipartite-maps-with-prescribed-
degrees/42BFC02A43167DDBBA5A3CC64F27BA35.

Izzo:2021:TGP

[75] Alexander J. Izzo and Dimitris Papathanasiou. Topology of Gleason parts in maximal ideal spaces with no analytic discs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(1):177–194, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/topology-of-gleason-parts-
in-maximal-ideal-spaces-with-no-
analytic-discs/15C9EDD01C869DC7387734C6DA93B2BD.

Pazzis:2021:PII

[76] Clément de Seguins Pazzis. Products of involutions of an infinitedimensional vector space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(1):195– 220, February 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/products-of-involutions-
                                        of-an-infinitedimensional-vector-
                                        space/CF40E13F1378C5DB01072729BBABB9EF.

Jiao:2021:MIN

[77] Yong Jiao, Fedor Sukochev, and Dejian Zhou. Maximal inequalities of noncommutative martingale transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (1):221–248, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/maximal-inequalities-of-
                                        noncommutative-martingale-transforms/
                                        48D3D4929C198BD022A4C1B2E04B9919.

Canakci:2021:EGA

[78] Ilke Çanakçi, David Pauksztello, and Sibylle Schroll. On extensions for gentle algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(1):249– 292, February 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-extensions-for-gentle-
                                        algebras/B8CF08D6367FF3773FC8DCFE028C2462.

Osaka:2021:EJA

[79] Hiroyuki Osaka and Tamotsu Teruya. Erratum: The Jiang–Su Absorption for Inclusions of Unital C∗-algebras.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (1):293–295, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/erratum-the-jiangsu-absorption-
for-inclusions-of-unital-calgebras/
5818143B89D6DF74DD853FD7C9E0075A.

See [1].

Anonymous:2021:CVIa

[80] Anonymous. CJM volume 73 issue 1 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (1):f1–f2, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-73-issue-1-cover-
and-front-matter/C4AEB73B7D5CD157AAB86CEBB6A562AF.

Anonymous:2021:CVIb

[81] Anonymous. CJM volume 73 issue 1 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (1):b1–b2, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-73-issue-1-cover-
and-back-matter/F6FED0F6F36C0A37D2139450D8347A13.

Karafyllia:2021:PHM

[82] Christina Karafyllia. On a property of harmonic measure on simply connected domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(2): 297–317, April 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-a-property-of-harmonic-
                                        measure-on-simply-connected-domains/
                                        8BC4364959434556539C972E8FB055C3.

Sendov:2021:SAG

[83] Blagovest Sendov and Hristo Sendov. Sector analogue of the Gauss–Lucas Theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(2):318–338, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/sector-analogue-
                                        of-the-gausslucas-theorem/E94D4CD87067D61234EF01833EBC2103.

Podesta:2021:NPS

[84] Fabio Podestà. Nearly parallel G_{2}- structures with large symmetry group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(2):339–359, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/nearly-parallel-g2structures-
                                        with-large-symmetry-group/E7F58AD649F3C1F89F0ABDFAC0C729E5.

Denisov:2021:GMR

[85] Sergey Denisov and Liban Mohamed. Generalizations of Menchov–Rademacher Theorem and existence of wave operators in Schrödinger evolution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (2):360–382, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/generalizations-of-menchovrademacher-
theorem-and-existence-of-wave-operators-
in-schrodinger-evolution/BD586B7845EA3CC0A6984D4E7C4FBAE7.

Ma:2021:LMC

[86] Dan Ma and Wei Wang. LYZ matrices and SL(n) contravariant valuations on polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(2): 383–398, April 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/lyz-matrices-and-sln-contravariant-
valuations-on-polytopes/0D4F5240F5825CED560B73B989AE6FBA.

Lemmens:2021:LOI

[87] Bas Lemmens, Onno van Gaans, and Hendrik van Imhoff. On the linearity of order-isomorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (2):399–416, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-linearity-of-orderisomorphisms/
C4EAB1A631B9D7D0BE5D55F7F71BA415.

Billig:2021:CSC

[88] Y. Billig and K. Iohara. Classification of simple cuspidal modules over a lattice Lie algebra of Witt type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(2):417–440, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/classification-of-simple-
cuspidal-modules-over-a-lattice-
lie-algebra-of-witt-type/611EDCFFEBD76E4329D3BFE5E440B9F9.

Nieto-Banos:2021:VMH

[89] Isidro Nieto-Baños and Pedro Luis del Angel-Rodriguez. Variation of mixed Hodge structures associated to an equisingular one-dimensional family of Calabi–Yau threefolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (2):441–464, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/variation-of-mixed-hodge-
structures-associated-to-an-equisingular-
onedimensional-family-of-calabiyau-
threefolds/062E8011A934FE330D53B36D1BE1DBA3.

Kellerhals:2021:IUP

[90] Ruth Kellerhals. Ideal uniform polyhedra in H^{n} and covolumes of higher dimensional modular groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (2):465–492, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/ideal-uniform-polyhedra-
in-mathbbhn-and-covolumes-of-higher-
dimensional-modular-groups/86B0CEF17F5E3146F719DE1022EA61B5.

Ueda:2021:MLP

[91] Yoshimichi Ueda. Matrix liberation process II: Relation to orbital free entropy. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(2):493–541, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/matrix-liberation-process-
ii-relation-to-orbital-free-entropy/
5909A9209DDF97B11A30D6CAFCAEA370.

Sakugawa:2021:GBE

[92] Kenji Sakugawa. Generalized Beilinson elements and generalized Soulé characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(2):542–571, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/generalized-beilinson-elements-
and-generalized-soule-characters/
B08F41641AF6FEE54CED590D02D5288B.

Monson:2021:UAS

[93] B. Monson and Egon Schulte. Universal alternating semiregular polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(2):572–595, April 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/universal-alternating-semiregular-
polytopes/EC048DE67D3BB3B0401F246509A6EB9D.

Anonymous:2021:CVIc

[94] Anonymous. CJM volume 73 issue 2 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (2):f1–f2, April 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-73-issue-2-cover-
and-front-matter/CA55E46BDFDFD5F853B6AE2060E17A6B.

Anonymous:2021:CVId

[95] Anonymous. CJM volume 73 issue 2 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (2):b1–b2, April 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-73-issue-2-cover-
and-back-matter/D3CEC688B3B0FD6233CEE48484F880BF.

Cheng:2021:CID

[96] Zhiyun Cheng. The chord index, its definitions, applications, and generalizations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3):597–621, June 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/chord-index-its-definitions-
applications-and-generalizations/
CB3FDAB7242B979CF8544AE66EC71662.

Chao:2021:BDT

[97] Zhang Chao and José L. Torrea. Boundedness of differential transforms for heat semigroups generated by Schrödinger operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3): 622–655, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/boundedness-of-differential-
transforms-for-heat-semigroups-
generated-by-schrodinger-operators/
F3532547EDCF38D1BF26D30D3630651A.

Fazelpour:2021:CDK

[98] Ziba Fazelpour and Alireza Nasr-Isfahani. Coxeter diagrams and the Köthe’s Problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3):656–686, June 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/coxeter-
diagrams-and-the-kothes-problem/
48D646BF31B040403EF9A9A84D416403.

Crooks:2021:ASG

[99] Peter Crooks and Maarten van Pruijssen. An application of spherical geometry to Hyperkähler slices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3): 687–716, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/an-application-of-spherical-
geometry-to-hyperkahler-slices/
BA55C84F689D171BD84EA7614BC626A3.

Kosinski:2021:EPU

[100] Lukasz Kosiński and Włodzimierz Zwonek. Extension property and universal sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3):717– 736, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/extension-property-and-
                                        universal-sets/F74578FAC08B97A34D124816F766982E.

Geudens:2021:CSS

[101] Stephane Geudens and Marco Zambon. Coisotropic submanifolds in b-symplectic geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3): 737–768, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/coisotropic-submanifolds-
                                        in-bsymplectic-geometry/62FF0471489FA8AF6E218A233DE9AC1A.

Iosevich:2021:REZ

[102] Alex Iosevich, Doowon Koh, Sujin Lee, Thang Pham, and Chun-Yen Shen. On restriction estimates for the zero radius sphere over finite fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3): 769–786, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-restriction-estimates-
                                        for-the-zero-radius-sphere-over-
                                        finite-fields/BD04A06192620C6B701E44B560F81F6A.

Kuwert:2021:RWS

[103] Ernst Kuwert and Tobias Lamm. Reflection of Willmore surfaces with free boundaries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3):787– 804, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/reflection-of-willmore-
surfaces-with-free-boundaries/392FC0A93E82D621F82ED81510C49E49.

Graham:2021:BSG

[104] Andrew Graham, Daniel R. Gulotta, and Yujie Xu. Bounding Selmer groups for the rankin-Selberg convolution of Coleman families. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3): 805–853, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/bounding-selmer-groups-
for-the-rankinselberg-convolution-
of-coleman-families/FF13C12CB84246804BE9509D89C6BCAB.

Shukla:2021:CSE

[105] Abhishek Kumar Shukla and Ben Williams. Classifying spaces for étale algebras with generators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3): 854–874, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/classifying-spaces-for-
etale-algebras-with-generators/
C65BFD4C7B056AE68A5A967658622E98.

Alsaody:2021:AAR

[106] Seidon Alsaody. Albert algebras over rings and related torsors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(3): 875–898, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/albert-algebras-over-rings-
                                        and-related-torsors/DFF3DF45F64477B98F71EC038A91428C.

Anonymous:2021:CVIe

[107] Anonymous. CJM volume 73 issue 3 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (3):f1–f2, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-73-issue-3-cover-
                                        and-front-matter/0C561EA225ED3DD078FFFFF80B4E553E.

Anonymous:2021:CVIf

[108] Anonymous. CJM volume 73 issue 3 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (3):b1–b2, June 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-73-issue-3-cover-
                                        and-back-matter/4368A6C6B726268BC628CCC042250D17.

Lewis:2021:FPC

[109] Joel Brewster Lewis and Alejandro H. Morales. Factorization problems in complex reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (4):899–946, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/factorization-problems-
in-complex-reflection-groups/5B3B3164BBD90F117D34F9D95094A7A9.

Bolanos:2021:TFC

[110] Wilmar Bolaños and Guillermo Mantilla-Soler. The trace form over cyclic number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(4):947–969, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/trace-form-over-cyclic-
number-fields/24F212C4CC6ACC8AB1527084B76F1999.

Anedda:2021:SSM

[111] Claudia Anedda, Fabrizio Cuccu, and Silvia Frassu. Steiner symmetry in the minimization of the first eigenvalue of a fractional eigenvalue problem with indefinite weight. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(4):970–992, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/steiner-
symmetry-in-the-minimization-of-
the-first-eigenvalue-of-a-fractional-
eigenvalue-problem-with-indefinite-
weight/F4E566F0A2581512E5A6652CC7C9881F.

Zhang:2021:BVL

[112] Mengyuan Zhang. Bundles on P^{n} with vanishing lower cohomologies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (4):993–1012, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/bundles-on-textbfpn-with-
vanishing-lower-cohomologies/822B90542051FBBE614FAE0D3BDE28FB.

Kajiwara:2021:DGS

[113] Tsuyoshi Kajiwara and Yasuo Watatani. Dimension groups for self-similar maps and matrix representations of the cores of the associated C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(4): 1013–1056, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/dimension-groups-for-selfsimilar-
maps-and-matrix-representations-
of-the-cores-of-the-associated-
calgebras/99ED5783021DD059A572F0BE531C0601.

Hu:2021:FAG

[114] Fei Hu and Sichen Li. Free abelian group actions on normal projective varieties: submaximal dynamical rank case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (4):1057–1073, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/free-abelian-group-actions-
on-normal-projective-varieties-
submaximal-dynamical-rank-case/
3F7DF4CFAF72600AB6C57431E966B28D.

Lemaire:2021:DEG

[115] Bertrand Lemaire and Jean-Loup Waldspurger. Données endoscopiques d’un groupe réductif connexe : applications d’une construction de Langlands. (French) [Endoscopic data of a related reductive group: applications of a Langlands construction]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(4): 1074–1094, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/donnees-endoscopiques-dun-
groupe-reductif-connexe-applications-
dune-construction-de-langlands/
42E52FAEE8D26B5F297CBABEFDBFC496.

Nandakumar:2021:CBM

[116] Vinoth Nandakumar and Gufang Zhao.

Categorification via blocks of modular representations for ∫↕_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(4): 1095–1123, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/categorification-via-blocks-
of-modular-representations-for-
mathfrak-sln/F9CCD325B631CDF9ABE74246A07046B5.

See addendum [179].

Marquis:2021:SKM

[117] Timothée Marquis. On the structure of Kac–Moody algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(4): 1124–1152, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/

                                        article/on-the-structure-of-kacmoody-
                                        algebras/ABF4DB7D5E80729588F11A96FCD4B142.

Roushon:2021:CSA

[118] S. K. Roushon. A certain structure of Artin groups and the isomorphism conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(4):1153– 1170, August 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/certain-structure-of-artin-
                                        groups-and-the-isomorphism-conjecture/
                                        1A6997A83B5F590AF674C910A1A8ED30.

See erratum [165].

Hua:2021:SRR

[119] Jiajie Hua and Qingyun Wang. Stability of rotation relations in C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (4):1171–1203, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/stability-of-rotation-relations-
                                        in-c-algebras/256E50605D81D875D779AE7391EBA602.

Anonymous:2021:CVIg

[120] Anonymous. CJM volume 73 issue 4 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(4): f1–f2, August 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/

article/cjm-volume-73-issue-4-cover-
and-front-matter/57C9130FB2D0079A889ACCAF43429539.

Anonymous:2021:CVIh

[121] Anonymous. CJM volume 73 issue 4 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (4):b1–b2, August 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-73-issue-4-cover-
and-back-matter/D61BC91A1453D2A7641C953DB45490BC.

Poveda:2021:SPF

[122] Alejandro Poveda, Assaf Rinot, and Dima Sinapova. Sigma–Prikry forcing I: The axioms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(5):1205– 1238, October 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/sigmaprikry-forcing-i-the-
axioms/BD09FDEAD0A7E4AEF5A557B207BAB121.

Willis:2021:KRH

[123] Michael Willis. Khovanov–Rozansky homology for infinite multicolored braids. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (5):1239–1277, October 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/khovanovrozansky-homology-
for-infinite-multicolored-braids/
D1A2DE52E85CEB10034ACD33E4CB519C.

Li:2021:SEE

[124] Hong-Quan Li and Peter Sjögren. Sharp endpoint estimates for some operators associated with the Laplacian with drift in Euclidean space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(5): 1278–1304, October 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/sharp-endpoint-estimates-
for-some-operators-associated-with-
the-laplacian-with-drift-in-euclidean-
space/B30CBE68F8222C406D1337C7DA0E75E9.

Kooistra:2021:TFM

[125] Remkes Kooistra and Alan Thompson. Threefolds fibred by mirror sextic double planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(5):1305– 1346, October 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/threefolds-fibred-by-mirror-
sextic-double-planes/3ECB07DB11FF9305DB2794AF3D76D244.

Magal:2021:VCF

[126] Pierre Magal and Ousmane Seydi. Variation of constants formula and exponential dichotomy for nonautonomous non-densely defined Cauchy problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (5):1347–1389, October 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/

article/variation-of-constants-
formula-and-exponential-dichotomy-
for-nonautonomous-nondensely-defined-
cauchy-problems/E822AC950D46E8A3C6166248A3D4C9DA.

Kleine:2021:BII

[127] Sören Kleine. Bounding the Iwasawa invariants of Selmer groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(5): 1390–1422, October 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/bounding-the-iwasawa-invariants-
of-selmer-groups/7C9DBEED5F8E5261F57F69A4F9D0E73D.

Daws:2021:RTF

[128] Matthew Daws and Bence Horváth.

Ring-theoretic (in)finiteness in reduced products of Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(5): 1423–1458, October 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/ringtheoretic-infiniteness-
in-reduced-products-of-banach-algebras/
23457CE0017B47F9B5B008364C69F0B1.

Anonymous:2021:CVIi

[129] Anonymous. CJM volume 73 issue 5 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(5): f1–f2, October 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/

                                        article/cjm-volume-73-issue-5-cover-
                                        and-front-matter/56E091CA6519ABD53458CBAF1D3C144E.

Anonymous:2021:CVIj

[130] Anonymous. CJM volume 73 issue 5 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (5):b1–b2, October 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-73-issue-5-cover-
                                        and-back-matter/44DF5631681222A59ACA4DB9B3516AFC.

Li:2021:LVD

[131] Junxian Li. Large values of Dirichlet L-functions at zeros of a class of L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(6):1459– 1505, December 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/large-values-of-dirichlet-
                                        lfunctions-at-zeros-of-a-class-
                                        of-lfunctions/C9F9C0BE6BBE62D41925ABE2B744291A.

Greither:2021:COC

[132] Cornelius Greither and Radan Kucera. On the compositum of orthogonal cyclic fields of the same odd prime degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (6):1506–1530, December 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-the-compositum-of-orthogonal-

cyclic-fields-of-the-same-odd-prime-
degree/1165F270D43F1AF619B6656FD07BEA54.

Angel:2021:MPS

[133] Omer Angel, Alexander E. Holroyd, Tom Hutchcroft, and Avi Levy. Mallows permutations as stable matchings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (6):1531–1555, December 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/mallows-permutations-as-
stable-matchings/491555EF74CDD241A225510C150EC89E.

Heittokangas:2021:NLI

[134] Janne Heittokangas, Hui Yu, and Mohamed Amine Zemirni. On the number of linearly independent admissible solutions to linear differential and linear difference equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(6): 1556–1591, December 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-number-of-linearly-
independent-admissible-solutions-
to-linear-differential-and-linear-
difference-equations/A89914ADB85632E0A772FA837B3CC845.

Goncalves:2021:EGA

[135] Daniel Gonçalves and Benjamin Steinberg. Étale groupoid algebras with coefficients in a sheaf and skew inverse semigroup rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(6):1592– 1626, December 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/etale-groupoid-algebras-
                                        with-coefficients-in-a-sheaf-and-
                                        skew-inverse-semigroup-rings/EBCD67C6E8CA938ACABB784A79855289.

Kabluchko:2021:NAW

[136] Zakhar Kabluchko, Daniel Temesvari, and Christoph Thäle. A new approach to weak convergence of random cones and polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(6): 1627–1647, December 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/new-approach-to-weak-convergence-
                                        of-random-cones-and-polytopes/68FCC081285821624BC9721FD3EC4954.

Furedi:2021:EPC

[137] Zoltán Füredi, Tao Jiang, Alexandr Kostochka, Dhruv Mubayi, and Jacques Verstraëte. Extremal problems for convex geometric hypergraphs and ordered hypergraphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(6):1648– 1666, December 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/extremal-problems-for-convex-
                                        geometric-hypergraphs-and-ordered-
                                        hypergraphs/E1CF12EB731D70BC2D5E1976693E3FD2.

Guo:2021:RCC

[138] Weichao Guo, Huoxiong Wu, and Dongyong Yang. A revisit on the compactness of commutators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(6): 1667–1697, December 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/revisit-on-the-compactness-
of-commutators/21B388C1ABA27409A0E3CC0B0EB3F86D.

Neumann:2021:KIB

[139] Daniel López Neumann. Kuperberg invariants for balanced sutured 3-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(6):1698– 1742, December 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/kuperberg-invariants-for-
balanced-sutured-3manifolds/D4420960263977933786C8314B8FFF71.

Massuyeau:2021:SFL

[140] Gwénaël Massuyeau and Delphine Moussard. A splicing formula for the LMO invariant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73(6):1743– 1770, December 2021. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/splicing-formula-for-the-
lmo-invariant/CFC2448E47E12465899F6A1123767264.

Anonymous:2021:CVIk

[141] Anonymous. CJM volume 73 issue 6 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (6):f1–f3, December 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-73-issue-6-cover-
                                        and-front-matter/0A074A35B97999DC45F965E580DEC28D.

Anonymous:2021:CVIl

[142] Anonymous. CJM volume 73 issue 6 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (6):b1–b2, December 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-73-issue-6-cover-
                                        and-back-matter/547BC8D4CE0B9BDD67378B5E00BBE116.

Arone:2022:SKC

[143] Gregory Arone and Markus Szymik. Spaces of knotted circles and exotic smooth structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(1):1–23, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/spaces-of-
                                        knotted-circles-and-exotic-smooth-
                                        structures/5B2F99D8401B2EA40B1515247F9D108A.

Farcaseanu:2022:FTC

[144] Maria Farcaseanu, Mihai Mihailescu, and Denisa Stancu-Dumitru. On a family of torsional creep problems in Finsler metrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(1):24– 41, February 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-a-family-of-torsional-
creep-problems-in-finsler-metrics/
ECFCE6955E427DD3EF91E1886DA41D38.

Lubbes:2022:MAS

[145] Niels Lubbes. Möbius automorphisms of surfaces with many circles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (1):42–71, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/mobius-automorphisms-of-
surfaces-with-many-circles/75B50FF6A827F24D37100571B40B9D77.

Backhausz:2022:ACO

[146] Ágnes Backhausz and Balázs Szegedy.

Action convergence of operators and graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(1):72–121, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/action-convergence-of-operators-
and-graphs/F5900CA5BE554C9F4DEAAB518962D2DD.

Gourevitch:2022:FCM

[147] Dmitry Gourevitch, Henrik P. A.

Gustafsson, Axel Kleinschmidt, Daniel Persson, and Siddhartha Sahi. Fourier coefficients of minimal and next-to-minimal automorphic representations of simply-laced groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(1):122– 169, February 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/fourier-coefficients-of-
                                        minimal-and-nexttominimal-automorphic-
                                        representations-of-simplylaced-
                                        groups/4A9BDB5DD9295F98A75929641CD52B72.

Airey:2022:SCR

[148] Dylan Airey, Steve Jackson, and Bill Mance. Some complexity results in the theory of normal numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (1):170–198, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/some-complexity-results-
                                        in-the-theory-of-normal-numbers/
                                        E81E3A923DA838FAB5699823A3939EA2.

Manon:2022:TGN

[149] Christopher Manon and Jihyeon Jessie Yang. Tropical geometry and Newton– Okounkov cones for Grassmannian of planes from compactifications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (1):199–231, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/tropical-geometry-and-newtonokounkov-
                                        cones-for-grassmannian-of-planes-
                                        from-compactifications/B51B04A4A587A19D15C91FFA600C97FF.

Lin:2022:DCT

[150] Jiawei Lin and Greg Martin. Densities in certain three-way prime number races. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(1):232–265, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/densities-in-certain-threeway-
prime-number-races/D70CBC7D8E8D158475E0D2A32D7EDA5F.

Cullinan:2022:DTS

[151] John Cullinan and Jeffrey Yelton. Divisibility of torsion subgroups of abelian surfaces over number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (1):266–298, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/divisibility-of-torsion-
subgroups-of-abelian-surfaces-over-
number-fields/B38F15B56C245BDABBBBE4197E700258.

Anonymous:2022:CVIa

[152] Anonymous. CJM volume 74 issue 1 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (1):f1–f3, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-74-issue-1-cover-
and-front-matter/0AC8A18A6F40A3BF5194E5EC127F6AE8.

Anonymous:2022:CVIb

[153] Anonymous. CJM volume 74 issue 1 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (1):b1–b2, February 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-74-issue-1-cover-
                                        and-back-matter/A12AC56CB6FBA5F266A4F1BAFB05A998.

Darby:2022:ECT

[154] Alastair Darby, Shintarô Kuroki, and Jongbaek Song. Equivariant cohomology of torus orbifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (2):299–328, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/equivariant-cohomology-
                                        of-torus-orbifolds/1E29F165087BA8837A7C04F14F3A0F16.

Lee:2022:SLM

[155] Man-Chun Lee and Luen-Fai Tam. Some local maximum principles along Ricci flows. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):329–348, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/some-local-maximum-principles-
                                        along-ricci-flows/7FF502B59589AD6AAB34339FE085BAA9.

Hoehn:2022:SPA

[156] L. C. Hoehn, L. G. Oversteegen, and E. D. Tymchatyn. Shortest paths in arbitrary plane domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):349–367, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/shortest-
paths-in-arbitrary-plane-domains/
748AF74BE7B99E99D1D6E940C2FA8849.

Lou:2022:FIP

[157] Yingying Lou, Dongmeng Xi, and Zhenbing Zeng. Further inequalities and properties of p -inner parallel bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):368–380, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/further-
inequalities-and-properties-of-
pinner-parallel-bodies/27DF41C510E9D4F62250EB8338B7F029.

Li:2022:PLA

[158] Qirui Li. A proof of the linear Arithmetic Fundamental Lemma for GL_{4}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):381–427, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/proof-of-
the-linear-arithmetic-fundamental-
lemma-for-operatorname-mathrm-gl4/
BC92E5548521AE74A5D564258ED1DB38.

Sargent:2022:OAO

[159] Meredith Sargent and Alan A. Sola.

Optimal approximants and orthogonal polynomials in several variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):428–456, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/optimal-approximants-and-
                                        orthogonal-polynomials-in-several-
                                        variables/D741863B935CC11B4E733B6535F71A96.

Jain:2022:DSE

[160] Tanvi Jain and Hemant Kumar Mishra. Derivatives of symplectic eigenvalues and a Lidskii type theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (2):457–485, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/derivatives-of-symplectic-
                                        eigenvalues-and-a-lidskii-type-
                                        theorem/FE3D8851687BCA387D8DF7E703F2602B.

Ye:2022:CCY

[161] Dongxi Ye. On a conjecture of Chen and Yui: Resultants and discriminants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):486–526, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-a-conjecture-of-chen-
                                        and-yui-resultants-and-discriminants/
                                        45C2DAA98564EB6C47F515EFE3465246.

Florens:2022:TIF

[162] Vincent Florens and Delphine Moussard. Torsions and intersection forms of 4-manifolds from trisection diagrams. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):527–549, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/torsions-
and-intersection-forms-of-4manifolds-
from-trisection-diagrams/48EE4B085494BE6143BC580E59EDF105.

Hambleton:2022:RCF

[163] Ian Hambleton and Semra Pamuk.

Rank conditions for finite group actions on 4-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):550– 572, April 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/rank-conditions-for-finite-
group-actions-on-4manifolds/D1C4A495F283217743832EBA2E7A87BB.

Gaiotto:2022:HSS

[164] Davide Gaiotto and Theo Johnson-Freyd. Holomorphic SCFTs with small index. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):573– 601, April 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/holomorphic-scfts-with-
small-index/8A7F04EAE2E9B59441584172CC294C25.

Roushon:2022:ECS

[165] S. K. Roushon. Erratum: A certain structure of Artin groups and the isomorphism conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(2):602, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/erratum-a-
                                        certain-structure-of-artin-groups-
                                        and-the-isomorphism-conjecture/
                                        23564BE22BE1E8F31A3775CD7F462860.

See [118].

Shimada:2022:CES

[166] Ichiro Shimada. Corrigendum: On an Enriques surface associated with a quartic Hessian surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (2):603–605, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/corrigendum-on-an-enriques-
                                        surface-associated-with-a-quartic-
                                        hessian-surface/88812D1442FA6F51029DAD8BF5BCFBCC.

See [2].

Anonymous:2022:CVIc

[167] Anonymous. CJM volume 74 issue 2 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (2):f1–f3, April 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-74-issue-2-cover-
                                        and-front-matter/829BAEA2336449955035DB2CBEA91909.

Anonymous:2022:CVId

[168] Anonymous. CJM volume 74 issue 2 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (2):b1–b2, April 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-74-issue-2-cover-
and-back-matter/3BBFDFF1D8D5855B1ED59F3E1D0A410B.

Mourrat:2022:PFH

[169] Jean-Christophe Mourrat. The Parisi formula is a Hamilton–Jacobi equation in Wasserstein space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3): 607–629, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/parisi-formula-is-a-hamiltonjacobi-
equation-in-wasserstein-space/ED0C76321ACC554426E1003BC96DB629.

Chandler:2022:TTR

[170] Alex Chandler, Adam M. Lowrance, Radmila Sazdanović, and Victor Summers. Torsion in thin regions of Khovanov homology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3):630– 654, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/torsion-in-thin-regions-
of-khovanov-homology/5B29A42168077C0DDA971F256CC6FD8B.

Eckhardt:2022:MGP

[171] Caleb Eckhardt, Kit Fieldhouse, Daniel Gent, Elizabeth Gillaspy, Ian Gonzales, and David Pask. Moves on kgraphs preserving Morita equivalence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3):655–685, June 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/moves-on-kgraphs-preserving-
                                        morita-equivalence/95D05FA1EBC8C40AF9181CEF4CE5D832.

Benkart:2022:MMF

[172] Georgia Benkart, Rekha Biswal, Ellen Kirkman, Van C. Nguyen, and Jieru Zhu. McKay matrices for finitedimensional Hopf algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3): 686–731, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/mckay-matrices-for-finitedimensional-
                                        hopf-algebras/7D171B78CC3EF96988C6E95CA5078147.

Bender:2022:RBP

[173] Chase Bender, Debraj Chakrabarti, Luke Edholm, and Meera Mainkar. L^{p}-regularity of the Bergman projection on quotient domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3): 732–772, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/lp-regularity-of-the-bergman-
                                        projection-on-quotient-domains/
                                        2E94772DE4F25DAAE0282243812DF34D.

Ostafe:2022:SPS

[174] Alina Ostafe and Igor E. Shparlinski. On the Skolem problem and some related questions for parametric families of linear recurrence sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3):773–792, June 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-skolem-problem-and-
some-related-questions-for-parametric-
families-of-linear-recurrence-sequences/
E5A153BF73722A2CE9C03B9D61CD9CF2.

Rose:2022:WQT

[175] David E. V. Rose and Logan C. Tatham.

On webs in quantum type C. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3): 793–832, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-webs-in-quantum-type-
c/61EB4E1E1D41A7EA73BA33F6BA9BED3E.

Dai:2022:DWI

[176] Feng Dai and Andriy Prymak. On directional Whitney inequality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3): 833–857, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-directional-whitney-
inequality/4D74D46DB2C8B76F20A06C1F5ED5EDAE.

Xue:2022:OI

[177] Hang Xue. Orbital integrals on GL_{n} × GL_{n} \ GL_{2n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3):858–886, June 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/orbital-
                                        integrals-on-textgln-times-textgln-
                                        backslash-textgl2n/47944D5E3DA533F071D7268DC5BB64A7.

Papageorgiou:2022:OMS

[178] Effie Papageorgiou. Oscillating multipliers on symmetric and locally symmetric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3):887– 905, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/oscillating-multipliers-
                                        on-symmetric-and-locally-symmetric-
                                        spaces/1CD992C1C81FF7CA690D382A93C5633A.

Nandakumar:2022:ACB

[179] Vinoth Nandakumar and Gufang Zhao. Addendum to: Categorification via blocks of modular representations for ∫↕_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(3):906, June 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/addendum-
                                        to-categorification-via-blocks-
                                        of-modular-representations-for-
                                        mathfrak-sln/3B31D516A92D57E22602C1C3C46AE7C1.

See [116].

Anonymous:2022:CVIe

[180] Anonymous. CJM volume 74 issue 3 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (3):f1–f3, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-74-issue-3-cover-
and-front-matter/734F167314A27CFF307109AC053FDB9B.

Anonymous:2022:CVIf

[181] Anonymous. CJM volume 74 issue 3 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (3):b1–b2, June 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-74-issue-3-cover-
and-back-matter/BCB04B365A572F98D47ED726E644FD33.

Budd:2022:SPP

[182] Timothy Budd and Nicolas Curien. Simple peeling of planar maps with application to site percolation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (4):907–941, August 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/simple-peeling-of-planar-
maps-with-application-to-site-percolation/
01C3AC8E90A0C34C6EF495623CB1A47F.

Fu:2022:TAS

[183] Xuanlong Fu and Huaxin Lin. Tracial approximation in simple C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (4):942–1004, August 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/tracial-approximation-in-
                                        simple-cast-algebras/94F6AD12DFFFF0A76B8B39DCB273B633.

Graczyk:2022:PKR

[184] P. Graczyk, T. Luks, and P. Sawyer. Potential kernels for radial Dunkl Laplacians. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(4):1005–1033, August 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/potential-kernels-for-radial-
                                        dunkl-laplacians/8D7683D1F242A6D45C4271DAD0A59FB0.

Biswas:2022:CTC

[185] Indranil Biswas, João Pedro dos Santos, Sorin Dumitrescu, and Sebastian Heller. On certain Tannakian categories of integrable connections over Kähler manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(4):1034–1061, August 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-certain-tannakian-categories-
                                        of-integrable-connections-over-
                                        kahler-manifolds/666803D13CDD812AB5332A1AD32D4710.

Fujita:2022:DDO

[186] Hajime Fujita. Deformation of Dirac operators along orbits and quantization of noncompact Hamiltonian torus manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(4):1062–1092, August 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/deformation-of-dirac-operators-
along-orbits-and-quantization-of-
noncompact-hamiltonian-torus-manifolds/
3D89134D79B7FCA1B1FCAAB26308D816.

Medvedev:2022:DSC

[187] Vladimir Medvedev. Degenerating sequences of conformal classes and the conformal Steklov spectrum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(4): 1093–1136, August 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/degenerating-sequences-
of-conformal-classes-and-the-conformal-
steklov-spectrum/596612496F19003D401FC542240FA714.

Friedl:2022:HRC

[188] Stefan Friedl, Takahiro Kitayama, Lukas Lewark, Matthias Nagel, and Mark Powell. Homotopy ribbon concordance, Blanchfield pairings, and twisted Alexander polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(4):1137– 1176, August 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/homotopy-ribbon-concordance-
blanchfield-pairings-and-twisted-
alexander-polynomials/6C8D79C8B75304C4C2794DC1461FDF10.

Agarwala:2022:CGWa

[189] Susama Agarwala, Siân Fryer, and Karen Yeats. Combinatorics of the geometry of Wilson loop diagrams I: equivalence classes via matroids and polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(4):1177– 1208, August 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/combinatorics-of-the-geometry-
                                        of-wilson-loop-diagrams-i-equivalence-
                                        classes-via-matroids-and-polytopes/
                                        6BA5E065F950C96E535719EF9DDFF994.

Anonymous:2022:CVIg

[190] Anonymous. CJM volume 74 issue 4 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(4): f1–f3, August 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-74-issue-4-cover-
                                        and-front-matter/ADCA815F1147107F5304BB2800FD066F.

Anonymous:2022:CVIh

[191] Anonymous. CJM volume 74 issue 4 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (4):b1–b2, August 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-74-issue-4-cover-
                                        and-back-matter/F0DA9D51ADE9622AFA9281882C535120.

Rego:2022:FLS

[192] Yuri Santos Rego. On the finiteness length of some soluble linear groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (5):1209–1243, October 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-finiteness-length-
of-some-soluble-linear-groups/666832220AA5B78FB5BF2B449EC4DEA2.

Kovac:2022:DTA

[193] Vjekoslav Kovac. Density theorems for anisotropic point configurations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5): 1244–1276, October 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/density-theorems-for-anisotropic-
point-configurations/E157E6A1EC3334C6BF9DF085C15FF532.

Akbal:2022:CEC

[194] Yildirim Akbal and Ahmet M. Güloglu.

Cyclicity of elliptic curves modulo primes in arithmetic progressions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5): 1277–1309, October 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cyclicity-of-elliptic-curves-
modulo-primes-in-arithmetic-progressions/
BD94A56E6424CFC71A636B80F5DA902C.

Hamaker:2022:IPD

[195] Zachary Hamaker, Eric Marberg, and Brendan Pawlowski. Involution pipe dreams. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5):1310– 1346, October 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/involution-pipe-dreams/
B6E8E918BC4BDB97543D20E2C70F4127.

Technau:2022:TCF

[196] Niclas Technau and Aled Walker. On the triple correlations of fractional parts of n^{2}α. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5):1347– 1384, October 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-triple-correlations-
of-fractional-parts-of-n2alpha/
DA519373102A5468FD50E22555A2AA7C.

Huynh:2022:TDS

[197] Tony Huynh and David R. Wood. Tree densities in sparse graph classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5): 1385–1404, October 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/tree-densities-in-sparse-
graph-classes/78244C7A924324B700B27A6EA08E1871.

Bley:2022:ECC

[198] Werner Bley and Alessandro Cobbe. The epsilon constant conjecture for higher dimensional unramified twists of Z_{p}^{r}(1). Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5):1405–1449, October 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (elec-

tronic). URL https://www.cambridge.
org/core/journals/canadian-journal-
of-mathematics/article/epsilon-
constant-conjecture-for-higher-
dimensional-unramified-twists-of-
mathbb-zpr-1/54762326E35D5D56E6811A1752A7E901.

Ye:2022:GTG

[199] Shuyang Ye. A group-theoretic generalization of the p-adic local monodromy theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5):1450– 1485, October 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/grouptheoretic-generalization-
of-the-padic-local-monodromy-theorem/
99D7B29FEE49CB9E4D8A7F514A8EBA62.

Xu:2022:PLC

[200] Ming Xu and Vladimir S. Matveev.

Proof of Laugwitz conjecture and Landsberg unicorn conjecture for Minkowski norms with SO(k) × SO(n − k)-symmetry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5):1486– 1516, October 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/proof-of-laugwitz-conjecture-
and-landsberg-unicorn-conjecture-
for-minkowski-norms-with-soktimes-
sonk-symmetry/5C065086CCAC3D4322824298006B41EC.

Anonymous:2022:CVIi

[201] Anonymous. CJM volume 74 issue 5 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(5): f1–f3, October 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-74-issue-5-cover-
                                        and-front-matter/3303D541561CE4C75276EAD87E76D147.

Anonymous:2022:CVIj

[202] Anonymous. CJM volume 74 issue 5 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (5):b1–b2, October 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-74-issue-5-cover-
                                        and-back-matter/6E64822C00907E8491721240E0CF4F42.

Parkinson:2022:AOS

[203] James Parkinson and Hendrik Van Maldeghem. Automorphisms and opposition in spherical buildings of exceptional type, I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(6):1517– 1578, December 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/automorphisms-and-opposition-
                                        in-spherical-buildings-of-exceptional-
                                        type-i/AD9065ECC91098A3DFD0366D7EBAC4D3.

Cibotaru:2022:NMS

[204] Daniel Cibotaru and Wanderley Pereira.

Nontame Morse–Smale flows and odd Chern–Weil theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(6):1579– 1624, December 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/nontame-morsesmale-flows-
and-odd-chernweil-theory/D4EC21C2A12686E85EEE5A399AC4625C.

Agarwala:2022:CGWb

[205] Susama Agarwala, Siân Fryer, and Karen Yeats. Combinatorics of the geometry of Wilson loop diagrams II: Grassmann necklaces, dimensions, and denominators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(6):1625– 1672, December 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/combinatorics-of-the-geometry-
of-wilson-loop-diagrams-ii-grassmann-
necklaces-dimensions-and-denominators/
2319B7070BAF3F1C98FC6CF898533C5C.

Ebrahimi-Fard:2022:WPN

[206] Kurusch Ebrahimi-Fard, Frédéric Patras, Nikolas Tapia, and Lorenzo Zambotti. Wick polynomials in noncommutative probability: a grouptheoretical approach. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(6):1673– 1699, December 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/wick-polynomials-in-noncommutative-
                                        probability-a-grouptheoretical-
                                        approach/BBD199A3C93DDDF7F7A99533A027FCD1.

Randrianantoanina:2022:IBN

[207] Narcisse Randrianantoanina. Interpolation between noncommutative martingale Hardy and BMO spaces: the case 0 < p < 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(6):1700– 1744, December 2022. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/interpolation-between-noncommutative-
                                        martingale-hardy-and-bmo-spaces-
                                        the-case-textbf-0p1/36E3744B7913D6D88F4EEED6A08AA29E.

Saksman:2022:AZH

[208] Eero Saksman and Odí Soler i Gibert. Approximation in the Zygmund and Hölder classes on R^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74(6): 1745–1770, December 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/approximation-in-the-zygmund-
                                        and-holder-classes-on-mathbb-rn/
                                        3FF38A152EEDD79EBFE79696D8F5B881.

Anonymous:2022:CVIk

[209] Anonymous. CJM volume 74 issue 6 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (6):f1–f3, December 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-74-issue-6-cover-
and-front-matter/B98BCF1911AF7CF15FF4226183E766CD.

Anonymous:2022:CVIl

[210] Anonymous. CJM volume 74 issue 6 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (6):f1–f3, December 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-74-issue-6-cover-
and-front-matter/B98BCF1911AF7CF15FF4226183E766CD.

Anonymous:2022:CVIm

[211] Anonymous. CJM volume 74 issue 6 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (6):b1–b2, December 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-74-issue-6-cover-
and-back-matter/2F5CECEE9D07C9A9959410AE73DF8D0F.

Anonymous:2022:CVIn

[212] Anonymous. CJM volume 74 issue 6 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (6):b1–b2, December 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/

article/cjm-volume-74-issue-6-cover-
and-back-matter/2F5CECEE9D07C9A9959410AE73DF8D0F.

Branding:2023:UCP

[213] Volker Branding, Stefano Montaldo, Cezar Oniciuc, and Andrea Ratto. Unique continuation properties for polyharmonic maps between Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(1):1–28, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/unique-continuation-
properties-for-polyharmonic-maps-
between-riemannian-manifolds/1785114B1C48398A7534BF8EDF5FD223.

Li:2023:AYU

[214] Wei Li, Alexey Ovchinnikov, Gleb Pogudin, and Thomas Scanlon. Algorithms yield upper bounds in differential algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(1):29– 51, February 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/algorithms-yield-upper-
bounds-in-differential-algebra/
5B0B3791CF7430ED1644CA1C6D5E2281.

Chepuri:2023:AKL

[215] Sunita Chepuri and Melissa Sherman-Bennett. 1324- and 2143-avoiding Kazhdan–Lusztig immanants and kpositivity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(1):52–82, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/1324-and-
2143avoiding-kazhdanlusztig-immanants-
and-kpositivity/12C0FD2A138C55B1F6800A82D6971D54.

Raghuram:2023:API

[216] A. Raghuram. An arithmetic property of intertwining operators for p-adic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):83–107, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/an-arithmetic-property-
of-intertwining-operators-for-padic-
groups/51F0259F85A91F4B6141DE076F38DEBA.

Langford:2023:PCP

[217] Jeffrey J. Langford. PDE comparison principles for Robin problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):108–139, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/pde-comparison-principles-
for-robin-problems/E4DC4553A4C970C63A8D786803F607D4.

Norton:2023:URT

[218] Emily Norton. Unitary representations of type B rational Cherednik algebras and crystal combinatorics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):140–169, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/

                                        canadian-journal-of-mathematics/
                                        article/unitary-representations-
                                        of-type-b-rational-cherednik-algebras-
                                        and-crystal-combinatorics/C14051F6F878960176F76D6AD9777A1B.

Hanusch:2023:CRP

[219] Maximilian Hanusch. A C^{k}–Seeleyextension-theorem for Bastiani’s differential calculus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(1):170– 201, February 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/mathcal-ck-seeleyextensiontheorem-
                                        for-bastianis-differential-calculus/
                                        9BD85B42863B240A3C7FA41A6FDA6073.

DiLorenzo:2023:CIR

[220] Andrea Di Lorenzo and Roberto Pirisi. Cohomological invariants of root stacks and admissible double coverings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):202–224, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cohomological-invariants-
                                        of-root-stacks-and-admissible-double-
                                        coverings/E1A9FB1DAB91E536471ABFF3D660FF9A.

Nicolau:2023:ITN

[221] Artur Nicolau and Pascal J. Thomas. Invertibility threshold for Nevanlinna quotient algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(1):225– 244, February 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/invertibility-threshold-
for-nevanlinna-quotient-algebras/
18FE2F79DB76A3EB897F5388A0906F10.

Polak:2023:PFR

[222] Elzbieta Polak and Dustin Ross.

Polynomiality of factorizations in reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(1):245– 266, February 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/polynomiality-of-factorizations-
in-reflection-groups/3989DAEAF96A61C2B04575FD30C1DAED.

Barmak:2023:IMG

[223] Jonathan Ariel Barmak. Invariants for metabelian groups of prime power exponent, colorings, and stairs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):267–297, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/invariants-for-metabelian-
groups-of-prime-power-exponent-
colorings-and-stairs/15BFD183D6CE9DC1FBC9AF09789AB56F.

Ray:2023:ECT

[224] Anwesh Ray and R. Sujatha. Euler characteristics and their congruences for multisigned Selmer groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):298–321, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/euler-characteristics-and-
                                        their-congruences-for-multisigned-
                                        selmer-groups/0B6062AEB078F7F238040DE2CB69A6A7.

Shkredov:2023:MES

[225] Ilya D. Shkredov. On multiplicative energy of subsets of varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):322–340, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-multiplicative-energy-
                                        of-subsets-of-varieties/DB81A73DCC86E2099016194B5FD5A095.

Anonymous:2023:CVIa

[226] Anonymous. CJM volume 75 issue 1 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):f1–f3, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-75-issue-1-cover-
                                        and-front-matter/DFA358877C9A42F1DE4FB74D14A85CCD.

Anonymous:2023:CVIb

[227] Anonymous. CJM volume 75 issue 1 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (1):b1–b2, February 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/

canadian-journal-of-mathematics/
article/cjm-volume-75-issue-1-cover-
and-back-matter/D3E8759AFB414C31F08E9E02A52526D7.

Bravi:2023:MSF

[228] Paolo Bravi and Jacopo Gandini. On the multiplication of spherical functions of reductive spherical pairs of type A. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):341–374, April 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-multiplication-of-
spherical-functions-of-reductive-
spherical-pairs-of-type-a/51479CA188DF6CA3C8EFE93A487E5FBA.

Asgari:2023:CAT

[229] Mahdi Asgari and Kiumars Kaveh. On combinatorics of the Arthur trace formula, convex polytopes, and toric varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):375–420, April 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-combinatorics-of-the-
arthur-trace-formula-convex-polytopes-
and-toric-varieties/308510E62F8AE8E1087C1635B0BEBC05.

Mathieu:2023:ESO

[230] Martin Mathieu and Michael Rosbotham. Exact structures for operator modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):421– 446, April 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/exact-structures-for-operator-
                                        modules/12F407ECFB85544F81DD1CEAA689BB11.

Aygin:2023:CGF

[231] Zafer Selcuk Aygin and Khoa D. Nguyen. N-colored generalized Frobenius partitions: generalized Kolitsch identities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):447–469, April 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/ncolored-
                                        generalized-frobenius-partitions-
                                        generalized-kolitsch-identities/
                                        6803C485644938F8E07CE72D70EE8CF2.

Cifre:2023:RPL

[232] María A. Hernández Cifre, Miriam Tárraga, and Jesús Yepes Nicolás. On the roots of polynomials with logconvex coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):470– 493, April 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-the-roots-of-polynomials-
                                        with-logconvex-coefficients/B727DED37F8DDE12613134126EEF0498.

Guzman:2023:UN

[233] Osvaldo Guzmán. The ultrafilter number and ⟨⇕. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):494– 530, April 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/ultrafilter-number-and-
mathfrak-hm/1C142C707040E1806368B566F02A5608.

Gaster:2023:CHM

[234] Jonah Gaster, Brice Loustau, and Léonard Monsaingeon. Computing harmonic maps between Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):531–580, April 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/computing-
harmonic-maps-between-riemannian-
manifolds/55AFB7A8E083007AE06824759234F8B4.

Gregory:2023:MTP

[235] Lorna Gregory. Maranda’s theorem for pure-injective modules and duality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):581–607, April 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/marandas-theorem-for-pureinjective-
modules-and-duality/37E826CCE7174A5167F8BE7F29FB4329.

Garrel:2023:MWR

[236] Nicolas Garrel. Mixed Witt rings of algebras with involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (2):608–644, April 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/

canadian-journal-of-mathematics/
article/mixed-witt-rings-of-algebras-
with-involution/4CB67233D5B5FA398969B6295A7FB29C.

Cavalcanti:2023:FSG

[237] Gil R. Cavalcanti, Ralph L. Klaasse, and Aldo Witte. Fibrations in semitoric and generalized complex geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(2):645–685, April 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/fibrations-in-semitoric-
and-generalized-complex-geometry/
5B0AFD74ADF43DEE8E9DCEBC75CD4745.

Anonymous:2023:CVIc

[238] Anonymous. CJM volume 75 issue 2 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (2):f1–f3, April 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-75-issue-2-cover-
and-front-matter/41BF283B27548A860C96B3A9B0F337F5.

Anonymous:2023:CVId

[239] Anonymous. CJM volume 75 issue 2 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (2):b1–b2, April 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-75-issue-2-cover-
and-back-matter/58E7D7A0934D57334CE8262269042E51.

Cojocaru:2023:BDF

[240] Alina Carmen Cojocaru and Tian Wang.

Bounds for the distribution of the Frobenius traces associated to products of non-CM elliptic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3): 687–712, June 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/bounds-for-the-distribution-
of-the-frobenius-traces-associated-
to-products-of-noncm-elliptic-curves/
5A5E87DF4896BEA1E69C75998620EAAC.

Khaochim:2023:ERA

[241] Chalinee Khaochim and Matthew A.

Papanikolas. Effective rigid analytic trivializations for Drinfeld modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3):713–742, June 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/effective-rigid-analytic-
trivializations-for-drinfeld-modules/
FA3D12981BE4291BB1A0213772C21773.

Genevois:2023:SCC

[242] Anthony Genevois. Special cube complexes revisited: a quasi-median generalization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3):743– 777, June 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/

                                        article/special-cube-complexes-
                                        revisited-a-quasimedian-generalization/
                                        18C2C6A5E6D97089D3A5E059A028CB16.

Franc:2023:EM

[243] Cameron Franc. Eisenstein metrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3):778–803, June 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/eisenstein-metrics/8D20202C419EE8B2CCA7046E52E632B2.

Chemotti:2023:GMS

[244] Frank Chemotti, Ján Minác, Andrew Schultz, and John Swallow. Galois module structure of square power classes for biquadratic extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3): 804–827, June 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/galois-module-structure-
                                        of-square-power-classes-for-biquadratic-
                                        extensions/D5D32774C956F93394D71D28C7505536.

Cao:2023:ESV

[245] Daomin Cao, Guolin Qin, and Changjun Zou. Existence of stationary vortex sheets for the 2D incompressible Euler equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3):828–853, June 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/existence-
                                        of-stationary-vortex-sheets-for-

the-2d-incompressible-euler-equation/
8DDF4525E95F0A97EE932B1F90B54438.

Ben-Neria:2023:CSR

[246] Omer Ben-Neria and Thomas Gilton.

Club stationary reflection and the special Aronszajn tree property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3): 854–911, June 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/club-stationary-reflection-
and-the-special-aronszajn-tree-
property/EFC83747E36349C9F23DE8BB817AF0CC.

Mori:2023:LTM

[247] Michiya Mori and Peter Semrl. Loewner’s theorem for maps on operator domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3):912–944, June 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/loewners-theorem-for-maps-
on-operator-domains/76488FE2CDABA4D1A4DEC047268D4E65.

Michaud-Jacobs:2023:ECP

[248] Philippe Michaud-Jacobs. On elliptic curves with p -isogenies over quadratic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3):945–964, June 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-elliptic-curves-with-
pisogenies-over-quadratic-fields/
7D814DC3E3A3EF4737B892AE975161D3.

Disegni:2023:AFL

[249] Daniel Disegni. p-adic L-functions via local-global interpolation: the case of GL_{2} × GU(1). Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(3):965– 1017, June 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/padic-lfunctions-via-localglobal-
                                        interpolation-the-case-of-mathrm-
                                        gl2times-mathrm-gutext1/F94EB647910E047F01FC642288C29ECD.

Mezer:2023:MOT

[250] Dor Mezer. Multiplicity one theorems over positive characteristic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (3):1018–1044, June 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/multiplicity-one-theorems-
                                        over-positive-characteristic/632149A0AB1674C2ED393DA4D827F97F.

Anonymous:2023:CVIe

[251] Anonymous. CJM volume 75 issue 3 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (3):f1–f3, June 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-75-issue-3-cover-
                                        and-front-matter/30D7EEF7B6E3AD0DD0CB88A58FF6D7CC.

Anonymous:2023:CVIf

[252] Anonymous. CJM volume 75 issue 3 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (3):b1–b2, June 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-75-issue-3-cover-
and-back-matter/8F4190BAF5D6DB5F8A178B4AE0C8EA76.

deJong:2023:TMT

[253] Robin de Jong and Farbod Shokrieh.

Tropical moments of tropical Jacobians. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (4):1045–1075, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/tropical-moments-of-tropical-
jacobians/7D168E767FECB100AF1FCF3EA6EF81F9.

Feigin:2023:TNG

[254] Evgeny Feigin, Martina Lanini, and Alexander Pütz. Totally nonnegative Grassmannians, Grassmann necklaces, and quiver Grassmannians. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4): 1076–1109, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/totally-nonnegative-grassmannians-
grassmann-necklaces-and-quiver-
grassmannians/F8B677DAE7631B365898112F96570702.

Bush:2023:SBC

[255] Michael Bush, Dale Frymark, and Constanze Liaw. Singular boundary conditions for Sturm–Liouville operators via perturbation theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4): 1110–1146, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/singular-boundary-conditions-
for-sturmliouville-operators-via-
perturbation-theory/A61B9B67836B3492038A515BAA6391A0.

Anderson:2023:GDG

[256] Theresa C. Anderson and Bingyang Hu. On the general dyadic grids onR^{d}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (4):1147–1175, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-general-dyadic-grids-
on-mathbb-rd/84AAE9718B3191E34483B52571323B05.

Fang:2023:TPR

[257] Xiang Fang and Pham Trong Tien. Two problems on random analytic functions in Fock spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4):1176– 1198, August 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/two-problems-on-random-
analytic-functions-in-fock-spaces/
4A56735D2DF825CDBA9794AB160217AE.

Chang:2023:QPG

[258] Wen Chang and Jie Zhang. Quivers with potentials for Grassmannian cluster algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4):1199– 1225, August 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/quivers-with-potentials-
for-grassmannian-cluster-algebras/
F28BDDDBF09D52D60EC14AC014BC4268.

Anjos:2023:LFG

[259] Sílvia Anjos, Miguel Barata, Martin Pinsonnault, and Ana Alexandra Reis. Loops in the fundamental group of Symp(CP^{2}♯5C̄P^{2}, ω) which are not represented by circle actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4): 1226–1271, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/loops-in-the-fundamental-
group-of-mboxsymp-mathbb-cmathbb-
p2-mbox5overline-mathbb-cmathbb-
p2omega-which-are-not-represented-
by-circle-actions/DA0F177C6A564C42EB7D7FA887ABB369.

Schotz:2023:GNT

[260] Matthias Schötz. Gelfand–Naimark theorems for ordered ∗-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4): 1272–1292, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/

                                        canadian-journal-of-mathematics/
                                        article/gelfandnaimark-theorems-
                                        for-ordered-algebras/7763CA0B82CA577DAAE672857E95D4A2.

Isik:2023:TDE

[261] Erman Isik, Yasemin Kara, and Ekin Ozman. On ternary Diophantine equations of signature (p, p, 3) over number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4):1293–1313, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-ternary-diophantine-
                                        equations-of-signature-pptext3-
                                        over-number-fields/A0B044F984B6B4B1968F5B126467BA3F.

Thiel:2023:GRS

[262] Hannes Thiel. The generator rank of subhomogeneous C^{∗} algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4): 1314–1342, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/generator-rank-of-subhomogeneous-
                                        calgebras/976DD3CC21BFD2F7688383FC08C8EE9C.

Bakic:2023:GGR

[263] Petar Bakić and Gordan Savin. The Gelfand–Graev representation of classical groups in terms of Hecke algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (4):1343–1368, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/

article/gelfandgraev-representation-
of-classical-groups-in-terms-of-
hecke-algebras/A49AC6D875350199DEB6EBE23C6DD6DC.

Pietrzycki:2023:TMC

[264] Paweł Pietrzycki and Jan Stochel. Twomoment characterization of spectral measures on the real line. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4): 1369–1392, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/twomoment-characterization-
of-spectral-measures-on-the-real-
line/688E707E017998F1C4970B43098B61F3.

Anonymous:2023:CVIg

[265] Anonymous. CJM volume 75 issue 4 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75(4): f1–f3, August 2023. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-75-issue-4-cover-
and-front-matter/23CA998312FA40FAA51C9CB6AEA31735.

Anonymous:2023:CVIh

[266] Anonymous. CJM volume 75 issue 4 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 75 (4):b1–b2, August 2023. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-75-issue-4-cover-
and-back-matter/44851A0C8D7DB33D72A41DCCD67755AD.